### v1

# Ruzsa’s problem on Bi-Sidon sets

Abstract.
A subset $S$ of real numbers is called bi-Sidon if it is a Sidon set with respect to both addition and multiplication, i.e., if all pairwise sums and all pairwise products of elements of $S$ are distinct. Imre Ruzsa asked the following question: What is the maximum number $f(N)$ such that every set $S$ of $N$ real numbers contains a bi-Sidon subset of size at least $f(N)$? He proved that $f(N)\geqslant cN^{1/3}$, for a constant $c>0$. In this note, we improve this bound to $cN^{5/12}$.
1. Introduction
A set of elements $S$ in an abelian group $(G,+)$ is called Sidon if for any elements $a,b,a^{\prime},b^{\prime}\in S$ we have $a+b=a^{\prime}+b^{\prime}$ if and only if $\{a,b\}=\{a^{\prime},b^{\prime}\}$. Sidon sets are classical objects in additive combinatorics and have been intensely studied over the decades [16, 20]. It was first proved by Singer [17] in 1938, that if $p$ is a power of a prime, then one can find $p+1$ integers such that their pairwise sums are all different modulo $p^{2}+p+1$. Since the ratio of consecutive primes tends to $1$, this immediately implies that $\{1,\ldots,N\}$ has an (additive) Sidon subset of size at least $(1-o(1))N^{1/2}$. Unaware of Singer’s result, three years later, Erdős and Turán [5] established a slightly weaker lower bound and an upper bound of $N^{1/2}+O(N^{1/4}).$ (For minor improvements, see [7, 2, 1].) Erdős repeatedly offered \$500 for a proof or disproof of the upper bound $N^{1/2}+o(N^{1/4}),$ but the problem is still open.
What happens if we replace the interval $\{1,\ldots,N\}$ with an arbitrary set $A\subset\mathbb{R}$ of size $N$? What is the size of the largest Sidon subset that can be found in any set $A$ of $N$ reals? Intuitively, the interval $\{1,\ldots,N\}$ has the richest “additive structure” among all sets of size $N$, so it should be the hardest to find Sidon subsets in it. This is indeed the case: Komlós, Sulyok and Szemerédi [6] proved that that any finite set $A\subset\mathbb{R}$ contains a Sidon subset $S$ of size at least $c|A|^{1/2}$. The simplest proof of this fact uses the so called “Ruzsa projection trick” [13], which allows us to “embed” any set $A$ into an arithmetic progression of comparable length while preserving all relevant linear relations between the elements in $A$. In this context, it does not really matter, whether we work over $\mathbb{Z},\mathbb{R},$ or $\mathbb{C}$. To make our note self-contained, in the next section, we include the short proofs of some well known facts that we need.
Similar statements hold for multiplicative Sidon sets $S\subset\mathbb{R}$, i.e., where all solutions of the equation $ab=a^{\prime}b^{\prime}$ in $S$ satisfy $\{a,b\}=\{a^{\prime},b^{\prime}\}$. We obtain that very subset $A\subset\mathbb{R}$ contains a multiplicative Sidon subset of size $c|A|^{1/2}$, for some $c>0$. The order of magnitude of this bound cannot be improved, as is shown by a geometric progression of length $N$.
More interesting things start to happen if we try to combine the additive and multiplicative structures of the real numbers. The exciting sum-product phenomenon in Additive Combinatorics postulates, roughly speaking, that finite subsets of real numbers cannot have very rich additive and multiplicative structures at the same time. For any $A\subset\mathbb{R}$, let $A+A=\{a+b\,|\;a,b\in A\},\;A\cdot A=\{a\cdot b\;|\;a,b\in A\}$. According to the celebrated Erdős–Szemerédi sum-product conjecture [4], for any finite $A\subset\mathbb{R}$, we have
$\max\{|A+A|,|A\cdot A|\}\geqslant|A|^{2-o(1)}$
as $|A|\rightarrow\infty.$ There has been a huge amount of work on the sum-product-type questions in many different settings, see, e.g., [3, 18, 12].
For any $A\subset\mathbb{R},$ let $S_{+}(A)$ and $S_{\times}(A)\subset A$ denote a maximum size additive Sidon subset and a maximum size multiplicative Sidon subset of $A$, respectively. We have
$\min(|S_{+}(A)|,|S_{\times}(A)|)\geqslant c|A|^{1/2},$
for a suitable constant $c>0$. Another, less known example of the sum-product phenomenon is a recent conjecture of Klurman and Pohoata [9], according to which any finite $A\subset\mathbb{R}$ satisfies
$\max(|S_{+}(A)|,|S_{\times}(A)|)\geqslant c|A|^{1/2+\varepsilon},$
for a suitable absolute constant $\varepsilon>0$. Roche-Newton and Warren [10] proved that if this conjecture is true, then we must have $\varepsilon\leqslant 1/6$.
In this note we consider a different way to combine addition and multiplication for Sidon sets, which was originated by Ruzsa [14]. We say that a set $S\subset\mathbb{R}$ is bi-Sidon if $S$ is both additive Sidon and multiplicative Sidon, i.e. for any $a,b,a^{\prime},b^{\prime}\in S$ the either of the equalities $a+b=a^{\prime}+b^{\prime}$ or $ab=a^{\prime}b^{\prime}$ implies that $\{a,b\}=\{a^{\prime},b^{\prime}\}$. Let $S_{bi}(A)$ denote a maximum size bi-Sidon subset of $A$. Ruzsa observed [15] that for $A\subset\mathbb{R}$, we have $|S_{bi}(A)|\gtrsim|A|^{1/3}$ and asked if this bound can be improved. Consider a couple of examples:
  • Let $A=\{1,\ldots,N\}$, then by the above we know that there is an additive Sidon subset $S_{+}\subset A$ of size at least $N^{1/2}$. On the other hand, the set $P$ of primes between $1$ and $N$ clearly forms a multiplicative Sidon subset in $A$ of size at least $(1+o(1))\frac{N}{\log N}$ by the Prime Number Theorem. So taking $S$ to be the intersection of $P$ and a random shift of $S_{+}$ we obtain a bi-Sidon subset in $\{1,\ldots,N\}$ of size at least $cN^{1/2}/\log N$.
  • Let $A=\{\gamma,\gamma^{2},\ldots,\gamma^{N}\}$ for an arbitrary $\gamma\neq-1,0,1$. Then we may consider $S_{\times}=\{\gamma^{s},~{}s\in S_{+}\}\subset A$ where $S_{+}\subset\{1,\ldots,N\}$ is the additive Sidon set from the previous example. Then $S_{\times}$ is a multiplicative Sidon subset in $A$ of size $N^{1/2}$. On the other hand, using some number theory, one can upper bound the number of quadruples $\gamma^{a}+\gamma^{b}=\gamma^{c}+\gamma^{d}$ and find a linearly sized additive Sidon subset $B\subset A$. Taking $S$ to be the intersection of $B$ and a random dilate of $S_{\times}$ produces a bi-Sidon subset of size $cN^{1/2}$.
These two extreme examples suggest that any set $A\subset\mathbb{R}$ should contain a bi-Sidon subset of size $|A|^{1/2+o(1)}$. In this note, we make the following improvement over the Ruzsa’s observation.
Theorem 1.
For any finite set $A\subset\mathbb{R}$ there exists a bi-Sidon subset $S\subset A$ of size at least $c|A|^{5/12}$.
Our proof is based on the energy sum-product estimate due to Rudnev–Shkredov–Stevens [11]. Their result states that any set $A\subset\mathbb{R}$ contains a subset $A^{\prime}$ of cardinality at least $|A|/2$ such that $A^{\prime}$ has small additive or multiplicative energy. We split into two cases depending on which energy is small. Suppose that $A^{\prime}$ has small multiplicative energy, the other case is analogous. We construct a random additive Sidon subset $B\subset A^{\prime}$ of size $\sim|A|^{1/2}$ which has the ‘expected’ multiplicative energy, i.e. each multiplicative quadruple in $A^{\prime}$ has the same probability of appearing in $B$ as if $B$ was a uniformly random subset in $A^{\prime}$ of size $|A|^{1/2}$. By taking a random subset of $B$ and deleting an element from each multiplicative quadruple we obtain a large bi-Sidon subset $S\subset A$. Here the gain is coming from the fact that $A^{\prime}$ had very few multiplicative quadruples which was then inherited by $B$ and so the deletion argument is more efficient. The construction of the sufficiently random Sidon subset $B$ is the crucial step of the argument and it proceeds by taking a uniformly random parabola $P\subset\mathbb{F}_{p}^{2}$ and lifting it to a subset of $A^{\prime}$ via an appropriate Freiman homomorphism.
A similar construction, based on a uniformly random parabola, was recently (and independently from this work) used by Tao [19] to provide a counterexample to an unrelated question of Erdős.
Our proof also works over other fields by using an appropriate energy sum-product estimate available over that field. For instance, using [11, Theorem 5] we get that if $A\subset\mathbb{F}_{p}$ has size at most $p^{5/8}$ then $A$ contains a bi-Sidon subset of size $c|A|^{2/5}$. More generally, for any $\varepsilon>0$ there exists $\delta>0$ such that if $A\subset\mathbb{F}_{p}$ has size at most $p^{1-\varepsilon}$, then $A$ contains a bi-Sidon subset of size at least $|A|^{1/3+\delta}$. On other hand, we do not know if there exists a bi-Sidon subset $S\subset\mathbb{F}_{p}$ of size at least $p^{1/3+\delta}$ for an absolute constant $\delta>0$.
Acknowledgements. We thank Imre Ruzsa for helpful conversations. The first author’s work was partially supported by ERC Advanced Grant “GeoScape”, no. 882971 and NKFIH (National Research, Development and Innovation Office) grant K-131529. The second author was supported by the Jane Street Graduate Fellowship.
2. Proof
Freiman homomorphisms. We will need a couple of standard results about Freiman homomorphisms. For the sake of completeness we include simple proofs.
Let $G,H$ be abelian groups and let $f:A\rightarrow H$ be a map defined on some subset $A\subset G$. We say that $f$ is a Freiman homomorphism on $A$ if for any elements $a,b,a^{\prime},b^{\prime}\in A$ such that $a+b=a^{\prime}+b^{\prime}$ it follows that $f(a)+f(b)=f(a^{\prime})+f(b^{\prime})$. A Freiman emdedding is a Freiman homomorphism $f:A\rightarrow H$ such that $f(a)\neq f(b)$ for any distinct $a\neq b\in A$ (i.e. $f$ is a Freiman isomorphism from $A$ to $f(A)$, see [20]).
Proposition 1.
Let $G$ be a torsion-free abelian group and $A\subset G$ is a finite subset. Then there exists a Freiman embedding $f:A\rightarrow\mathbb{Z}$.
Proof.
We may assume that $G\cong\mathbb{Z}^{d}$. Let $n_{1}<n_{2}<\ldots<n_{d}$ be a sequence of integers and define for $a=(a_{1},\ldots,a_{d})\in A$:
$f(a)=n_{1}a_{1}+n_{2}a_{2}+\ldots+n_{d}a_{d}\in\mathbb{Z}.$
Clearly $f$ is a Freiman homomorphism and it is an embedding if we choose $n_{1},\ldots,n_{d}$ to be growing fast enough.    $\Box$
Proposition 2.
Let $A\subset\mathbb{Z}$ be a set of size $N$. For any prime $p\geqslant 2$ and and an integer $d\geqslant 1$ such that $N\leqslant 2^{-2d-1}p^{d}$, there exists a subset $A^{\prime}\subset A$ of size at least $2^{-d-1}N$ and a Freiman embedding $f:A^{\prime}\rightarrow\mathbb{F}_{p}^{d}$.
Proof.
Define $\varphi_{p}:\mathbb{R}\rightarrow\mathbb{F}_{p}$ to be the map $\varphi_{p}(x)=[px]\pmod{p}$. Let $\theta_{1},\ldots,\theta_{d}\in[0,1]$ be uniformly random and independent numbers and define a (random) map $f:\mathbb{Z}\rightarrow\mathbb{F}_{p}^{d}$ by
$f(a)=(\varphi_{p}(\theta_{1}a),\ldots,\varphi_{p}(\theta_{d}a)).$
Let $B\subset A$ be the subset of elements $a\in A$ such that $\{p\theta_{i}a\}\in[0,1/2)$ for every $i=1,\ldots,d$ (here $\{x\}$ denotes the fractional part of a real number $x$). Observe that $f$ restricted to $B$ is a Freiman homomorphism: indeed, for each $i=1,\ldots,d$ observe that
$\varphi_{p}(\theta_{i}b)+\varphi_{p}(\theta_{i}b^{\prime})=[p\theta_{1}b]+[p\theta_{i}b^{\prime}]=[p\theta_{i}(b+b^{\prime})]=\varphi_{p}(\theta_{i}(b+b^{\prime})),$
for $b,b^{\prime}\in B$. Observe also that $\mathbb{E}|B|=2^{-d}N$ since $\theta_{i}$ are independent and the fractional part $\{n\theta_{i}\}$ is distributed uniformly in $[0,1)$ for any integer $n$.
Let $C\subset A$ be the set of elements $a\in A$ such that there exists $a^{\prime}\neq a$ with $f(a)=f(a^{\prime})$. Note that for fixed distinct elements $a\neq a^{\prime}\in A$ we have
$\Pr[f(a)=f(a^{\prime})]=\Pr[\varphi_{p}(\theta_{1}a)=\varphi_{p}(\theta_{1}a^{\prime})]^{d}\leqslant\Pr[\operatorname{dist}(\theta_{1}(a-a^{\prime}),\mathbb{Z})\leqslant 1/p]^{d}\leqslant(2/p)^{d}.$
So by the union bound, the probability that $a\in C$ is bounded by $N(2/p)^{d}\leqslant 2^{-d-1}$. We conclude that $\mathbb{E}|C|\leqslant 2^{-d-1}N$. Finally, we observe that if we let $A^{\prime}=B\setminus C$ then $f:A^{\prime}\rightarrow\mathbb{F}_{p}^{d}$ is a Freiman embedding and $\mathbb{E}|A^{\prime}|\geqslant\mathbb{E}|B|-\mathbb{E}|C|\geqslant 2^{-d-1}N$. So there exists a choice of $\theta_{1},\ldots,\theta_{d}$ which achieves $|A^{\prime}|\geqslant 2^{-d-1}N$.    $\Box$
Proof of Theorem 1. Let $A\subset\mathbb{R}$ be a finite set. It is safe to assume that $A\subset\mathbb{R}_{>0}$. By the energy sum-product estimate due to Rudnev–Shkredov–Stevens [11, Theorem 5], there exists an absolute constant $C>0$ and a subset $A^{\prime}\subset A$ of size at least $|A|/2$ such that
(1) $\min\{E^{+}(A^{\prime}),E^{\times}(A^{\prime})\}\leqslant C|A|^{3-1/4},$
where $E^{+}(A^{\prime})$ and $E^{\times}(A^{\prime})$ denote additive and multiplicative energies of $A^{\prime}$, respectively, defined by
$E^{+}(A^{\prime})=\{(a,b,a^{\prime},b^{\prime})\in A^{\prime 4}:~{}a+b=a^{\prime}+b^{\prime}\},$
$E^{\times}(A^{\prime})=\{(a,b,a^{\prime},b^{\prime})\in A^{\prime 4}:~{}ab=a^{\prime}b^{\prime}\}.$
We therefore split into two cases depending on which energy in (1) is smaller. Suppose that $E^{\times}(A^{\prime})\leqslant C|A|^{3-1/4}$ holds. Let $p$ be a prime number such that $p\in[8|A|^{1/2},16|A|^{1/2}]$. By combining Propositions 1 and 2 we can find a subset $A^{\prime\prime}\subset A^{\prime}$ of size at least $|A^{\prime}|/8$ and a Freiman embedding $f:A^{\prime\prime}\rightarrow\mathbb{F}_{p}^{2}$.
Let $P\subset\mathbb{F}_{p}^{2}$ be a uniformly random parabola, namely, let $P_{0}=\{(t,t^{2}),~{}t\in\mathbb{F}_{p}\}$ be the standard unit parabola, take $g\in\operatorname{Aff}(\mathbb{F}_{p}^{2})$ to be a uniformly random invertible affine map $g:\mathbb{F}_{p}^{2}\rightarrow\mathbb{F}_{p}^{2}$ and define $P=g(P_{0})$.
Observation 3.
Let $v_{1},v_{2},v_{3},v_{4}\in\mathbb{F}_{p}^{2}$ be distinct vectors. Then we have $\Pr[\{v_{1},v_{2},v_{3}\}\subset P]\leqslant p^{-3}$ and $\Pr[\{v_{1},v_{2},v_{3},v_{4}\}\subset P]\leqslant 4p^{-4}$.
Proof.
If some three of the vectors $v_{i}$ are collinear then they cannot lie on a parabola and so the probability in question is 0 in this case. So we may assume that no three of $v_{i}$-s are collinear.
Note that the probability $q=\Pr[\{v_{1},v_{2},v_{3}\}\subset P]$ does not depend on the choice of non-collinear vectors $v_{1},v_{2},v_{3}$. Indeed, any 2 non-collinear triples $v_{1},v_{2},v_{3}$ and $u_{1},u_{2},u_{3}$ can be mapped to each other by some affine map $h$ and since the distribution of $P$ is $h$-invariant, the probabilities are the same. Since there are $p^{2}(p^{2}-1)(p^{2}-p)$ non-collinear tuples $(v_{1},v_{2},v_{3})$ in $\mathbb{F}_{p}^{2}$ and every parabola $P$ contains exactly $p(p-1)(p-2)$ of them, we conclude by a double-counting argument that
$q=\frac{p(p-1)(p-2)}{p^{2}(p^{2}-1)(p^{2}-p)}=\frac{p-2}{p^{2}(p+1)(p-1)}.$
Thus, it remains to prove the upper bound the conditional probability
$q^{\prime}=\Pr[v_{4}\in P~{}|~{}v_{1},v_{2},v_{3}\in P].$
The probability above is invariant under affine transformations of vectors $v_{i}$ so, by applying an appropriate affine map, we may assume that $v_{1}=(0,0)$, $v_{2}=(1,0)$, $v_{3}=(0,1)$ and $v_{4}=(x_{0},y_{0})$ for some $x_{0},y_{0}\neq 0$ and $x_{0}+y_{0}\neq 1$ (here we use that no 3 vectors among $v_{i}$ are collinear). Note that there are exactly $p-2$ parabolas $P$ which pass through the points $(0,0)$, $(1,0)$ and $(0,1)$ and that every point $(x_{0},y_{0})$ belongs to at most 2 of them. One way to see this is to write down a parameterization of this family of parabolas in coordinates explicitly. Namely, for $s\in\mathbb{F}_{p}\setminus\{0,1\}$ put
$P_{s}=\{(x,y)\in\mathbb{F}_{p}^{2}:~{}(x+sy)^{2}=x+s^{2}y\}.$
This implies that $q^{\prime}\leqslant\frac{2}{p-2}$ and so $qq^{\prime}\leqslant\frac{2}{p^{2}(p^{2}-1)}\leqslant 4p^{-4}$.    $\Box$
Now define $B=f^{-1}(P)$, i.e. the set of elements $a\in A^{\prime\prime}$ which are getting mapped to the random parabola $P$. Since $P\subset\mathbb{F}_{p}^{2}$ is a Sidon set, $B$ is an additive Sidon subset in $A$. Clearly, $\mathbb{E}|B|=p^{-1}|A^{\prime\prime}|\geqslant 2^{-7}|A|^{1/2}$. We can also estimate the multiplicative energy of $B$. For convenience, let $E_{0}^{\times}(B)$ denote the number of multiplicative quadruples $b_{1}b_{2}=b_{3}b_{4}$ with all $b_{i}$ distinct and $E_{1}^{\times}(B)$ the number of quadruples $b_{1}b_{2}=b_{3}b_{4}$ with $b_{1}=b_{2}$ or $b_{3}=b_{4}$ but not both.
Let $a_{1}a_{2}=a_{3}a_{4}$ be a multiplicative quadruple in $A^{\prime\prime}$, then
  • If the elements $a_{1},a_{2},a_{3},a_{4}$ are all distinct then by Observation 3 applied to vectors $v_{i}=f(a_{i})$, we obtain $\Pr[\{a_{1},a_{2},a_{3},a_{4}\}\subset B]\leqslant 4p^{-4}$,
  • If $a_{1}\neq a_{2}$ and $a_{3}=a_{4}$, then by Observation 3 we obtain $\Pr[\{a_{1},a_{2},a_{3},a_{4}\}\subset B]\leqslant p^{-3}$. Similarly for the case $a_{1}=a_{2}$ and $a_{3}\neq a_{4}$.
We conclude by linearity of expectation
$\mathbb{E}[E_{0}^{\times}(B)]\leqslant 4p^{-4}E^{\times}(A^{\prime\prime})\leqslant 4Cp^{-4}|A|^{3-1/4}\leqslant C|A|^{3/4},$
$\mathbb{E}[E_{1}^{\times}(B)]\leqslant p^{-3}|A^{\prime\prime}|^{2}\leqslant|A|^{1/2}.$
Let $q=c|A|^{-1/12}$ for some small enough $c>0$ and let $\tilde{B}\subset B$ be a random subset where each element is included independently with probability $q$. We then have
$\mathbb{E}|\tilde{B}|=q\mathbb{E}|B|\geqslant 2^{-7}c|A|^{5/12},$
$\mathbb{E}[E_{0}^{\times}(\tilde{B})]=q^{4}\mathbb{E}[E_{0}^{\times}(B)]\leqslant c^{4}C|A|^{5/12},$
$\mathbb{E}[E_{1}^{\times}(\tilde{B})]=q^{3}\mathbb{E}[E_{1}^{\times}(B)]\leqslant c^{3}|A|^{1/4}.$
Finally, let $S\subset B$ be a set obtained from $B$ by removing an element from each non-trivial multiplicative quadruple. It is clear that then $S$ is a bi-Sidon set and its expected size is at least
$\mathbb{E}|S|\geqslant\mathbb{E}|\tilde{B}|-\mathbb{E}[E_{0}^{\times}(\tilde{B})]-\mathbb{E}[E_{1}^{\times}(\tilde{B})]\geqslant(2^{-7}c-c^{4}C-c^{3})|A|^{5/12}$
so taking $c>0$ small enough and choosing $S$ with $|S|\geqslant\mathbb{E}|S|$ gives us the desired bi-Sidon subset $S\subset A$. All of this was done in the case when the second term achieves the minimum in (1). In the complementary case, we have the bound $E^{+}(A^{\prime})\leqslant C|A|^{3-1/4}$ instead. It is easy to see that exactly the same argument applies and so we get a bi-Sidon subset $S\subset A^{\prime}$ of size $c^{\prime}|A|^{5/12}$ in this case too.